! Required information The Foundational 15 (Algo) (LO5-1, LO5-3, LO5-4, LO5-5, LO5-6, LO5-7, LO5-8] [The following informatio6. If the selling price increases by $2 per unit and the sales volume decreases by 100 units, what would be the net operating7. If the variable cost per unit increases by $1, spending on advertising increases by $1,400, and unit sales increase by 1808. What is the break-even point in unit sales? Break-even point units9. What is the break-even point in dollar sales? Break-even point10. How many units must be sold to achieve a target profit of $13,500? Number of units
Answer #1
Part 6 Oslo Company
Contribution Income Statement
Amount Per unit % of Sales
Sales(900 units) $                       46,800 $                   52.00 100.00%
Variable Costs $                       24,750 $                   27.50 52.88%
Contribution Margin $                       22,050 $                   24.50 47.12%
Fixed Costs $                       14,850
Net Operating Income $                         7,200
Part 7 Oslo Company
Contribution Income Statement
Amount Per unit % of Sales
Sales(1,180 units) $                       59,000 $                   50.00 100.00%
Variable Costs($27.50+$1.00) $                       33,630 $                   28.50 57.00%
Contribution Margin $                       25,370 $                   21.50 43.00%
Fixed Costs($14,850+$1,400) $                       16,250
Net Operating Income $                         9,120
Part 8 Break-even Point as per the original Income Statement =$14,850 / $22.50 =660 units
Break-even Point as per Part 7 =$14,850 / $24.50 =606 units
Break-even Point as per Part 8 =$16,250 / $21.50 =756 units
Part 9 Break-even Point as per the original Income Statement =$14,850 / 45% =$33,000
Break-even Point as per Part 7 =$14,850 / 47.1115384615% =$31,518
Break-even Point as per Part 8 =$16,250 / 43% =$37,791
Part 10 As per Original Income Statement
Target Profit $                       13,500
Add:Fixed expenses $                       14,850
Target Contribution margin $                       28,350
Contribution Margin per unit $                         22.50
No. of units required 1260 units
As per Part 6 Income Statement
Target Profit $                       13,500
Add:Fixed expenses $                       14,850
Target Contribution margin $                       28,350
Contribution Margin per unit $                         24.50
No. of units required 1157 units
As per Part 7 Income Statement
Target Profit $                       13,500
Add:Fixed expenses $                       16,250
Target Contribution margin $                       29,750
Contribution Margin per unit $                         21.50
No. of units required 1384 units
Note:For Part 8,9,10 it has not been mentioned that which Income statement is to be considered, Therefore I solved all three